# AURA: A Hybrid Spatiotemporal-Chromatic Framework for Robust, Real-Time Detection of Industrial Smoke Emissions

### v1

@@ Abstract
This paper introduces AURA, a novel hybrid spatiotemporal-chromatic framework designed for robust, real-time detection and classification of industrial smoke emissions. The framework addresses critical limitations of current monitoring systems, which often lack the specificity to distinguish smoke types and struggle with environmental variability. AURA leverages both the dynamic movement patterns and the distinct color characteristics of industrial smoke to provide enhanced accuracy and reduced false positives. This framework aims to significantly improve environmental compliance, operational safety, and public health outcomes by enabling precise, automated monitoring of industrial emissions.
@@ 4.4.1 Spatiotemporal Noise-Cancelled Motion Score
[…]
@@ 4.4.2 Multi-Point Chromatic Aberration Metric (Delta E)
[…]
@@
Where $S_{C}$ is the chromatic score, $\mathbf{C}(i,j,t)$ is the pixel’s CIELAB vector, $\bar{\mathbf{C}}_{bg}(t)$ is the dominant background profile, and k-NN()($\cdot$)( ⋅ ) is the clustering operator.
@@ 6.3 Analysis of Results
Faster RCNN 0.721
YOLOv8s 0.881
Fire and Smoke detection model [6] 0.878
AURA (Ours) 0.988

### v2

@@ Abstract
This paper introduces AURA (A Hybrid Spatiotemporal-Chromatic Framework), a novel system for robust, real-time detection and classification of industrial smoke emissions. Conventional monitoring systems and standard deep learning models struggle with the challenges of industrial environments, including distant, low-resolution targets, significant camera instability, and the difficulty of distinguishing smoke from atmospheric phenomena like steam and fog. AURA overcomes these limitations through a dual-engine architecture. A spatiotemporal engine isolates true plume motion using an innovative noise cancellation technique that decouples it from camera shake, while a parallel chromatic engine analyzes spectral deviations in the perceptually uniform CIELAB color space. This hybrid approach not only enhances detection accuracy but also enables actionable classification of smoke by type (e.g., black, white), addressing a critical gap in existing systems. Validated on an extensive 1,100-hour dataset of real-world industrial video, AURA achieved a 98.80% F1-score, significantly outperforming conventional models and demonstrating remarkable resilience under severe environmental conditions. The framework provides a scalable and effective solution for enhancing environmental compliance, operational safety, and industrial process control.
@@ 3.4.1 Spatiotemporal Noise-Cancelled Motion Score
To isolate true plume motion from environmental noise, we define a Spatiotemporal Noise-Cancelled Motion Score, $S_{M}$, for each pixel $(i,j)$ at time $t$. This score is calculated by subtracting a dynamically generated motion noise profile, $\Psi$, from the raw motion energy, $\mathcal{M}$, derived from an adaptively selected background subtraction model, $\mathcal{B}_{k}$. A Heaviside step function, $\Theta$, nullifies the score when global scene instability, $\sigma_{G}$, exceeds a threshold, $\tau_{shake}$.
@@ 3.4.2 Multi-Point Chromatic Aberration Metric (Delta E)
The chromatic aberration score, $S_{C}$, for a pixel $(i,j)$ is the perceptual color distance (CIEDE2000, denoted $\Delta E_{00}$) between the pixel’s CIELAB value, $\mathbf{C}(i,j,t)$, and a dynamically computed dominant background chrominance profile, $\bar{\mathbf{C}}_{bg}(t)$. This profile is the centroid of the largest cluster from a k-NN analysis on multiple background reference points, $\{\mathbf{R}_{n}\}$.
@@
Where $S_{C}$ is the chromatic score, $\mathbf{C}(i,j,t)$ is the pixel’s CIELAB vector, $\bar{\mathbf{C}}_{bg}(t)$ is the dominant background profile, and k-NN()($\cdot$) is the clustering operator.
@@ 5.3 Analysis of Results
Faster RCNN 72.1
YOLOv8s 88.1
Fire and Smoke detection model [6] 87.8
AURA (Ours) 98.8